### v1

# Assessing the Impact of Upselling in Online Fantasy Sports

Abstract.
This study explores the impact of upselling on user engagement. , we model users’ deposit behaviour on the fantasy sports platform Dream11. Subsequently, we develop an experimental framework to evaluate the effect of upselling using an intensity parameter. Our live experiments on user deposit behaviour reveal decreased user recall with heightened upselling intensity. Our findings indicate that increased upselling intensity improves user deposit metrics and concurrently diminishes user satisfaction and conversion rates. We conduct robust counterfactual analysis and train causal meta-learners to personalise user’s upselling intensity levels to reach an optimal trade-off point.
Behavioural sciences, Online Upselling, Causal Machine Learning, Price Sensitivity, Meta-learners
journalyear: 2024copyright: acmlicensedconference: The Fourth International Conference on Artificial Intelligence and Machine Learning Systems; ; Louisiana State University, Baton Rouge, LAbooktitle: The Fourth International Conference on Artificial Intelligence and Machine Learning Systems (AIMLSystems 2024), October 25–28, 2024, Bangalore, Indiaisbn: 979-8-4007-1649-2/23/10
1. Introduction
Upselling or cross-selling refers to offering users additional products or services that are higher-end or complementary to their initial purchase. This strategy is widely employed across various industries to enhance revenue and customer value. According to Schiffman (Schiffman, 2005), ”Upselling is the initiative to encourage customers who have already purchased to buy more of the same product or additional products.”, whereas Kubiak et al. (2010) describe upselling as advancing to a more expensive product or service (Kubiak and Weichbroth, 2010).
Upselling is studied extensively across multiple sectors, such as airlines, hospitality, retail, and insurance. And their impacts on customer satisfaction and revenue metrics. For instance, in the airline industry, upselling involves flight seat upgrades (Thirumuruganathan et al., 2023), while in the hospitality industry, it involves luxury room upgrades (Guillet, 2020),(Guelman et al., 2014) Insurance industry to cross-sell/Upsell.
In this research paper, we study user behaviour after the Dream11 business’s upselling and develop a system to optimise the upselling policy, which has not been effectively addressed by existing methods relying on traditional machine learning techniques.
In this work, we draw the following conclusions to the field of user behaviour modelling in upselling on online platforms:
  • Using a multi-class supervised model, we predict the deposit propensity of a user and evaluate it across supervised regression, classification and heuristic baseline methods. During online experiments, we analysed the trade-offs between recall and revenue uplift.
  • Evaluate causal models to address the trade-offs and arrive at an optimal upselling policy from real-world deposit data collected using grid experiments across the spectrum of intensity values.
By accurately predicting deposit amounts and tailoring upselling suggestions, we aim to ensure a seamless payment system, user engagement, fairness, and the platform’s revenue optimisation. We conducted several online grid experiments across intensity gradients to upsell users. We found that increasing intensity leads to a decrease in user recall. The recall dropped because some users disliked upselling, and many new users had dropped off the conversion funnel. Therefore, it was evident that there should be a personalised upselling policy. The proposed approach builds upon previous work in uplift modelling to estimate heterogeneous treatment effects (Zhao and Harinen, 2019).
2. Problem Formulation and Algorithms
We develop a supervised model to predict the expected deposit amount for each user. This model will feed the upselling strategy by generating a list of suggested deposit amounts tailored to individual users.
2.1. Supervised Model Formulation
Let $\mathbf{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ denote the set of user-specific and deposit-specific features, including demographic information, historical transaction data, and engagement metrics. Let $F(\mathbf{X})$ represent the set of deposit amounts for a user, modelled as a multi-class classification model. The problem involves two key components: predicting the expected deposit amount for each user and computing the recommended list of suggested deposit amounts. We want to learn the function $F(\mathbf{X})$ from the historical deposit data $\mathbf{X}$ to predict the expected deposit amounts.
2.1.1. Evaluation and Learnings
In this paper, we introduce key metrics designed to provide insights into the accuracy and reliability of our classification models. Since our dataset was imbalanced towards lower deposit amounts, we selected the weighted F1 score as our evaluation criterion because it is crucial to have a model that is equally correct for all the classes. The dataset is imbalanced with lower-value transaction amounts, causing a significant load on the payment systems. Focal loss and l2 regularization proved pivotal for the accuracy of the final model.
Model F1 Score (Weighted)
Heuristic (Median rolling 10 Txn) 0.736
LGBM Regressor 0.758
LGBM Classifier 0.804
LGBM Classifier (Focal Loss) 0.852
Table 1. Model Performance Comparison
3. Experimentation
3.1. Upselling Experiment
The payments dataset used in this research comes from millions of payment transactions conducted on the Dream11 fantasy sports gaming platform. Whenever the user adds cash to the account, we recommend a set of three values shown in table 2. There are maximum limits and UX constraints on these values. We arrived at these values by analyzing historical deposit amounts and identifying the most frequent deposit amounts and their multiples. This allowed us to cover a range of common deposit behaviours and encourage higher deposits without overwhelming the user. Below are
Amount to add
Prefill value Option 1 Option 2
50 200 500
Table 2. Recommended Deposit Values
The objective was to experiment with upselling by recommending higher deposit amounts with guardrails on the following user engagement metrics. Deposit funnel: Total users attempted to make a transaction -¿ Total users successfully made a transaction. User Conversion: Paid registrations /Total registrations Using an intensity factor ($a$) and the user deposit propensity values, we implemented the following grid experimentation strategy:
  1. (1)
    Sample deposit amounts from $\text{F}(X_{i})$ for each user.
  2. (2)
    Multiply intensity ($a$) to the prediction and arrive at the recommended values.
For each user ($i$), we randomly assigned the intensity treatments to 4 different target groups (TGs), as given in Table [3].
Target Group Treatment
TG1 1x
TG2 1.25x
TG3 1.5x
TG4 2x
Table 3. Intensity Treatment Groups
3.2. Online Results
The control group received a default recommendation, whereas the target groups received various personalized amounts alongside different discrete intensity values [1x, 1.25x, 1.5x, 2x]. This experiment lasted eight weeks, from January 1, 2024, to February 28, 2024. We observed an immediate impact on the per transaction value of about +5.6%, while users were also making fewer transactions on the application, resulting in a 4.8% fall in the most aggressive variant. While evaluating the long-term sustained impact in the avg_deposit_week_8, Figure [1] illustrates the uplift across the experimental variants shown in Table [3].
Figure 1. Recall and Revenue Uplift Across Variants
Upselling impacted the experiment guardrails, which were the user engagement metrics. The time spent for a successful payment increased, and user conversion dropped by almost 2.1%, mainly impacting new users.
4. Causal Uplift Modelling
While analyzing the experiment of varying the intensity parameter, The challenge was to estimate how much to upsell a user without a trade-off in user satisfaction. A popular approach for efficiently assigning treatments to a subset of customers for targeted personalization is using causal machine learning models (Gutierrez and Gérardy, 2017; Diemert, 2018). For any causal uplift modelling task, we need to gather data on three essential items: Treatment (T), Response (Y), and Feature / Covariate (X).
4.1. Dataset Preparation
We split the data into two parts: data_train and data_test collected from the randomized control trials in Section 3
  • Treatment (Ti): Intensity [1.0, 1.25, 1.50, 2.0]—
  • Response Metric (Y): Deposit
  • Xi: User Features, Gameplay Features, User Demographics
We can arrive at an optimal upselling policy by personalising the intensity factor for each user, We estimate the incremental uplift after the intensity treatments (Ti) is given to each user uniformly. Next, we will use this dataset to model upselling.
4.2. Model Training
We estimated the CATE $Y(x)$, the conditional increment in the probability of completing the deposit $\Pr(Y=1)$ due to upselling intensity, given user features (X). We tested the following algorithms: S-learner, T-learner, X-learner, and R-learner, implemented in Python using (Chen et al., 2020), with XGBRegressor as the base learner and a propensity score of 0.5.
For training, we used the hyperopt package (Bergstra et al., 2013) with SparkTrials for distributing our workloads and managed multiple model runs with the MLflow library (Zaharia et al., 2018). We used the Tree-based Parzen Estimators (TPE) for efficient parameter search and defined the objective function as follows:
  • Initiate Models: Obtain models from causalml (Chen et al., 2020) and define the parameter search space.
  • Train Models on Training Data: Train the models using the training dataset.
  • Infer Models: Apply the trained models to the training and test datasets.
  • Create Validation/Optimization Metric: Calculate the mean of the test set AUUC (Area Under the Uplift Curve) scores for validation and optimization.
We conducted 1000 random trials, logging the AUUC for each treatment on the test set. We evaluated two model configurations: The global configuration, selecting the model with the lowest test set loss, and the local configuration, predicting the Conditional Average Treatment Effect (CATE) for the i-th treatment using the model with the highest test set i-th AUUC.
Table [4] summarises the configurations for the XGBRegressor used in the spark trials. Each treatment group (TG1, TG2, TG3, TG4) and the global model (TG_global) have tailored local configuration parameters, while the global configuration is optimized for overall performance, and local configurations are optimised for individual treatments. We compare the CATE scores for each treatment and keep a threshold of greater than zero else a control treatment is assigned to the user as shown below.
$\pi(x)=\begin{cases}\operatorname{argmax}\limits_{t\neq\text{CG}}\text{CATE}(x\mid T=t)&\text{if}\;\max\limits_{t\neq\text{CG}}\text{CATE}(x\mid T=t)>\text{0}\\ \text{CG}&\text{otherwise}\end{cases}$
Parameter TG1 TG2 TG3 TG4 TG_global meta_learner r_learner x_learner r_learner s_learner r_learner gamma 1.24 8.40 7.95 4.60 5.67 n_estimators 59 153 102 125 200 colsample_bytree 0.82 0.56 0.93 0.88 0.75 max_depth 7 10 15 8 12 min_child_weight 5 3 6 4 2 reg_lambda 0.36 0.14 0.57 0.94 0.50 reg_alpha 77 115 50 163 90 Configuration Local config Global config
Table 4. XGBRegressor Configurations
4.3. Evaluation
Uplift modelling faces performance limitations due to the need for an unbiased evaluation metric. Individual responses to actions and natural responses are unknown simultaneously, preventing explicit labelling of uplift responses. We underwent rigorous evaluations on both the business and model sides to establish confidence in our analysis.
  • Percent Treated: Percentage of users distributed across all possible treatment values, including control. In Table [5], the grid-experiment policy is uniformly distributed while the CATE policy is assigned with a non-negative threshold.
    Percent Treated policy
    Treatment Grid exp policy CATE policy
    CG 19.92 3.16
    TG1 20.12 12.71
    TG2 20.05 31.86
    TG3 20.13 37.42
    TG4 19.78 14.86
    Table 5. Policy Assignment Table
  • ERUPT: Expected Response Under Proposed Treatments: (Zhao et al., 2017), (Hitsch and Misra, 2018) It quantifies a model’s performance at the downstream task of predicting the optimal treatment to assign each member of the population. It is defined as follows
    (1) $\text{ERUPT}=\mathbb{E}(Y\cdot\mathbb{I}(\pi(x)=t)\mid T=t,\pi)$
    We calculate the expected bootstrapped revenue with 5000 as the sample size. The CATE policies Test(pi) and Train(pi) arrive after model inference, and the intensity factor with the highest uplift is selected. For the CATE policy, Figure [2] shows a significant 10.7% uplift in both train and test datasets.
    Figure 2. Erupt revenue distribution for grid experiment policy vs CATE policy across test and train datasets.
  • AUUC There is an Uplift Curve and a random curve based on the calculation of every treatment. The Uplift charts indicate that there is good offline fitment of the underlying estimators, Cum gain above is plotted by ordering the users based on the inferred uplift scores, and we observe that most of the gains are captured in the initial set of users.
    Figure 3. AUUC curves across treatments, the y-axis represents the cumulative incremental gains, and the x-axis is the proportion of the population targeted.
5. Conclusion and Future Work
5.1. Conclusion
Our research demonstrates the complex trade-offs between upselling intensity and user satisfaction. We observed that increasing intensity in upselling strategies led to higher immediate revenue but decreased user recall and satisfaction. This was particularly evident in the TG1 group, where the bi-modal distribution of user responses indicates a polarized reaction to upselling efforts. These findings underscore the importance of balancing upselling intensity with user experience to ensure user engagement and revenue uplift.
The causal uplift modelling framework helps us determine a personalised intensity value for each user, balancing the trade-offs and ensuring fairness in the recommendations. The causal analysis results across both test and train datasets indicate that taking the models online can be promising, with an expected revenue uplift of 10.7%.
We further analysed the online results in Figure [1], the guardrails were majorly affected for new users in the system. The offline CATE policy treats with 1x intensity to 80% of new users, leading to an impressive 2.1% improvement in conversion rates. While the offline results are promising, we must still validate the optimal trade-offs in an online experimental setting.
5.2. Future Work
Our future work will focus on testing the models online, exploring additional user features, and extending the experimentation to incentivise user behaviour upselling on the platform. Further, developing personalized upselling strategies that dynamically adjust intensity levels based on user behaviour and feedback could enhance revenue and user satisfaction.
6. Appendix
6.1. Platform Implementation
The ”Deposit Amount Upselling” platform shown in figure 4 integrates several key components to predict deposit amounts and provide personalized upselling recommendations. Below is a summary of the implementation steps:
  • Data Ingestion: The platform ingests data from two primary sources: streaming events and backend data stored in the warehouse.
  • Feature Aggregation/pre-processing: We prepare the data by performing the following steps: label creation, gradient features calculation, outlier removal and feed it to the Deposit Amount Predictor model, ensuring the model leverages both real-time and historical data for accurate predictions.
  • Deposit Amount Predictor: A machine learning model predicts the likely deposit amounts for users. The model outputs deposit amount classes (e.g., 0-20, 20-40, 1000-25000) and corresponding probability scores.
  • Upselling Recommendation: The predicted amounts and scores are processed by an upselling policy, which employs an intensity layer to experiment using the model’s predictions.
6.2. Technical Stack
  • Data Ingestion: Real-time data processing is performed using Kafka on our in-house stream processing platform (Gaikwad, 2023). Backend data is retrieved from SQL databases like Redshift.
  • Feature Store: Processed features are stored efficiently in a feature store, supporting high throughput and low-latency retrieval. Delta Lake is used for offline storage, while Redis is employed for online storage.
  • Model Training and Prediction: Causal Machine learning models are developed in Python using (Chen et al., 2020) library. Airflow manages the scheduling for both batch and real-time inference scenarios.
  • Model Management: ML Flow (Zaharia et al., 2018) provides seamless model tracking and registry management.
  • Deployment and Inference: We power and monitor our inference jobs on our in-house machine-learning platform, Darwin
  • Backend Integration: Recommendations are served into backend systems for delivery to end-user devices.
Figure 4. Personalised Upselling Model